OpenStudy (anonymous):

How many real number solutions are there to the equation 0 = –3x² + x – 4?

OpenStudy (anonymous):

Hi, do you know how to solve quadratic equations?

OpenStudy (anonymous):

I haven't learned this, but yet it's on my study quide, Idk why...

OpenStudy (anonymous):

That's OK, no problem. So you don't know how to solve this, right: \[x ^{2}+5x+6=0\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

No, I don't know how to do any of thiss.

OpenStudy (anonymous):

OK. Should I teach you how, or do you just want the answer to your original question?

OpenStudy (anonymous):

Any help would be great.

OpenStudy (anonymous):

OK, the answer to your question is 2. I can teach you how to solve these, if you want?

OpenStudy (anonymous):

Thanks!---Yeah, if you have the time and know how to teach people tht really don't understand half this stuff lol

OpenStudy (zarkon):

\[–3x² + x – 4=-3\left(x-\frac{1}{6}\right)^2-\frac{47}{12}\] this is always negative and thus never zero...so no real solutions

OpenStudy (anonymous):

OK, sure! Before I go on, though, I made a mistake, the answer should have been 0, sorry :P

OpenStudy (anonymous):

Ohh, okay!

OpenStudy (anonymous):

multiply by -1 on both sides. =3x^2-x+4 use quadratic equation to find value of x x=0.1667 + 1.1426i x= 0.1667 - 1.1426i this will give you complex root .There is not real root.

OpenStudy (anonymous):

Right. The easiest way to solve these is what's called "Completing the Square". I was going to type out a little lecture, but maybe it's easier to show you this video: http://www.youtube.com/watch?v=gzm-uhj06q8 It's 13 minutes long, and hopefully by the end you'll understand it all. Probably much easier to understand than any written explanations from me or anybody.

OpenStudy (anonymous):

quadratic equation is just a equation, your teacher may give you prove of this in next lectures. if you don't know this then that easy take ur equation that is -3x2+x-4=0 exponent 2 coefficeint is a=-3 exponent 1 coefficient is b=1 exponent 0 coefficeint is c=-4 just put these in quadratic equation that is this.

OpenStudy (anonymous):

Thanks you guys!!

OpenStudy (anonymous):

No problem :D

OpenStudy (anonymous):

visit this link that will you good stuff about math. http://www.mathsisfun.com/algebra/quadratic-equation.html